Your accounts receivable clerk, Mitra Adams, to whom you pay a salary of $2,535 per month, has just purchased a new Acura. You decide to test the accuracy of the accounts receivable balance of $138,580 as shown in the ledger. The following information is available for your first year in business. (1) Collections from customers $334,620 (2) Merchandise purchased 540,800 (3) Ending merchandise inventory 152,100 (4) Goods are marked to sell at 40% above cost Compute an estimate of the ending balance of accounts receivable from customers that should appear in the ledger and any apparent shortages. Assume that all sales are made on account.

The ending balance of accounts receivable from customers $

Apparent shortage $
Answer #1

Cost of gods sold = Merchandise purchased -  Ending merchandise inventory

= 540,800 - 152,100 = $388,700

Sales = Cost of gods sold + Cost of gods sold*40%

= $388,700 + $388,700*0.40 = $544,180

The ending balance of accounts receivable from customers = Sales - Collections from customers

= $544,180 - $334,620

= $209,560

Apparent shortage = $209,560 -  $138,580 = $70,980